# Models Matter: Setting Accurate Privacy Expectations for Local and Central Differential Privacy

### v1

@@ Abstract.
differential privacy, privacy-enhancing technologies, usable privacy
ccs: Security and privacy Usability in security and privacyccs: Human-centered computing Human computer interaction (HCI)ccs: Security and privacy Human and societal aspects of security and privacy
1. Introduction
@@
Usable privacy research has long focused on making information about privacy protections transparent to end users in order to allow for informed decision-making about data sharing (Lipford et al., 2009; Felt et al., 2016; Emami-Naeini et al., 2022). Prior work has sought to explain privacy enhancing technologies (PETs) such as end-to-end encryption in messaging (Demjaha et al., 2018) and HTTPS/TLS (Felt et al., 2014). Existing explanation strategies vary in terms of how much they try to explain data protection mechanisms or processes versus implications of the particular PET. As PETs become increasingly complex and offer more nuanced notions of privacy, there is a significant need for increased research into best practice for transparent PET messaging that matches these new techniques and empowers end users.
@@
Differential privacy (DP) (Dwork et al., 2006a) is a privacy-enhancing technology that has quickly been adopted by industry and government agencies (Erlingsson et al., 2014; Messing et al., 2020; Thakurta et al., 2017; Abowd, 2018; Miklau, 2021; Desfontaines, 2021). DP deployments provide provable privacy guarantees by adding statistical noise to data or computations; this noise obfuscates the information of each individual while preserving aggregate-level insights. In response to DP’s rapid success, a growing body of work has started to document the inadequacies of existing messaging around DP (Cummings et al., 2021) and design new messaging techniques for DP systems (Xiong et al., 2020, 2022; Karegar et al., 2022; Nanayakkara et al., 2023; Bullek et al., 2017; Smart et al., 2022; Franzen et al., 2022).
@@
The technical and mathematical complexity of PETs like DP makes effective communication challenging (Cummings et al., 2021; Nanayakkara and Hullman, 2022; Abu-Salma et al., 2018; Demjaha et al., 2018). DP is an interesting case study for constructing transparent PET messaging because it is an instance of an emerging PET paradigm that has received relatively little attention: privacy-preserving, outsourced computation. This paradigm is increasingly important as more users rely on computationally-weak mobile devices (Pew Research Center, 2021). In response to this growing need, new approaches to privacy-preserving computational outsourcing are being actively developed, both in industry (Amazon Web Services, [n. d.]; Microsoft, [n. d.]) and academia (Lapets et al., 2018). Although there has been some work on people’s mental models of other PETs in this category (Kacsmar et al., 2022) and creating transparent messaging for functional encryption in particular (Alaqra et al., 2023), further work is needed. DP is particularly interesting because of its widespread deployment and use of statistical methods (as opposed to encryption).
@@
In this work, we develop messaging for DP that highlights the threat models that are implicit in different approaches to deploying DP. These implicit threat models are critical for end users to understand before sharing their data, as the chosen threat model may not provide protections against the classes of attackers about which they are concerned. We do this by exploring three explanation formats drawn from the existing PETs messaging literature: nutrition labels (Kelley et al., 2009), diagrams (Xiong et al., 2022; Smart et al., 2022; Stransky et al., 2021), and metaphors (Zhang-Kennedy et al., 2013; Demjaha et al., 2018; Raja et al., 2011; Karegar et al., 2022; Zhang-Kennedy et al., 2016; Suh et al., 2022). Each of our evaluated explanations aims to communicate the consequences of DP in terms of which information flows the deployment protects against. Prior work has tended to focus on explaining individual PETs in isolation, and while we focus on DP, we discuss how our designs may be modified to account for the effects of multiple PETs deployed together.
@@
Differential Privacy Deployment Models. There are multiple deployment models for DP, each of which is associated with a particular threat model. The two most widely deployed models are the central model (Dwork et al., 2006b) and the local model (Kasiviswanathan et al., 2011). [^1] Although there are many variations of DP (Desfontaines and Pejó, 2020), we focus on the central and local models due to their popularity.
@@
The central model assumes there exists a data curator who is trusted to see raw data from individuals; the adversary can only access released, aggregate results. The data curator collects data from individuals, performs statistical analyses on the collected dataset, and then injects statistical noise into the results before release. This process limits the ability of the adversary to reconstruct individual records from summary statistics, at the cost of reduced accuracy. The potential danger of this model, however, is that the data curator might not be trustworthy; the database storing individuals’ data could be vulnerable to hackers or misuse by insiders if other, complementary security practices are not adopted in tandem. Well-known deployments of the central model include the U.S. Census Bureau’s data products for the 2020 Decennial Census (Abowd, 2018).
@@
In the local model, noise is added to each individual’s data before collection, meaning the unmodified data are never stored together. As such, there is not the same need to trust the data curator (i.e., it is assumed that the data curator is honest but curious). This higher level of security comes at a cost: significantly more noise must be added to the data in order to ensure the same level of privacy protection, reducing the accuracy—and, thus, utility—of the collected data. Notably, Google and Apple have both used local DP to analyze browser data in Chrome and Safari, respectively (Erlingsson et al., 2014; Apple, 2017).
@@
Data subjects cannot be expected to make informed data-sharing decisions if they believe that DP is “some sort of crypto-magic to protect people from data misuse” (Rogaway, 2015). Prior work has demonstrated that existing DP description strategies do a poor job aligning users’ privacy expectations with the privacy protections provided by DP models (Cummings et al., 2021). In other words, the kind of protection that users expect does not align with the actual nature of the protection offered by DP. Misaligned expectations exacerbated by poor communication can lead to data subjects underestimating or—even more alarmingly—overestimating the privacy protection that DP offers.
@@
We first explore three kinds of explanations that build on prior work  (Karegar et al., 2022; Near and Darais, 2020; Xiong et al., 2022; Kelley et al., 2009; Raja et al., 2011; Cummings et al., 2021) through an interview study: metaphors, diagrams, and privacy labels for DP. Based on the results of this study, we identify the most promising strategies—privacy labels and metaphors—and further refine these explanations based on participant feedback. We evaluate our refined explanations in an online survey ($n=698$), measuring objective comprehension, subjective understanding, perceived thoroughness, and trust. We compare our explanations against existing state-of-the-art text-based explanations of DP (Xiong et al., 2020). Based on our results, we make suggestions for future research and design of explanations of PETs.
@@
2. Background
A growing body of work provides guidance for effective S&P communication (Schaub et al., 2015; Gluck et al., 2016a). Awkward interfaces or ineffective communication can lead to dangerous misconceptions and risky behaviors (Whitten and Tygar, 1999; Krombholz et al., 2019; Cummings et al., 2021; Habib et al., 2018). One reason that people may misjudge privacy risks or misuse PETs is that they lack appropriate mental models. Prior work has argued that “efficacy of risk communication depends not only on the nature of the risk, but also on the alignment between the conceptual model embedded in the risk communication and the user’s mental model of the risk” (Asgharpour et al., 2007). Unfortunately, existing depictions of DP appear to be misaligned with people’s mental models, resulting in misaligned privacy expectations (Cummings et al., 2021).
@@ 2. Background
Implications vs Process. One of the most important findings from prior work is that explaining data protection processes is not enough for most readers to grasp the implications of the protection offered by PETs (Smart et al., 2022; Cummings et al., 2021; Xiong et al., 2020; Distler et al., 2020). Xiong et al. (Xiong et al., 2020) studied explanations of both central and local DP, and found that when the implications of the local and central models were stated explicitly, participants were more willing to share information under the local model. Kühtreiber et al. (Kühtreiber et al., 2022) replicated this study with German participants. Differently from these studies, we explore a variety of best-practice methods from the usable S&P literature (i.e., metaphors, diagrams, and nutrition labels) to communicate which information flows are protected by DP under the central and local models.
Table 1. Five Information Disclosures. We combine the “organization” and “data analyst” categories from prior work, since a data analyst is simply an employee of the organization (Cummings et al., 2021). Although some implementations of the central model limit employees’ access to the data (e.g., Uber (Johnson et al., 2020)), we consider the more common case where only published information is privacy-protected.
@@
Cummings et al. (Cummings et al., 2021) explored the implications of DP through six information disclosures about which people care and DP may protect—depending on whether the local or central model is used. They found that existing descriptions of DP fail to appropriately set privacy expectations regarding these disclosures, in part because many descriptions are not specific to the model (e.g., central or local) being used. In contrast to this work, we build new explanations rather than evaluating existing ones. We draw on their framework to present the implications of DP (entities to whom data can potentially be disclosed) as part of our nutrition labels. For improved clarity, in our study, we combine two of the disclosures (organization and data analyst), resulting in five total (Table 1).
@@
Finally, Frazen et al. (Franzen et al., 2022) and Nanayakkara et al. (Nanayakkara et al., 2023) developed methods of explaining the implications of the privacy budget, drawing from the risk communication literature. Nanayakkara et al. (Nanayakkara et al., 2023) found that participants were more willing to share information as the privacy budget decreased (i.e., protections were strengthened). In our study, we assume a small privacy budget (i.e., strong privacy), so that we can focus on the implications of the deployment model. Future work could consider combining our explanations with explanations of the privacy budget.
@@
Metaphors Metaphors are one approach for improving mental models, and have been studied extensively in the S&P domain (Zhang-Kennedy et al., 2013; Demjaha et al., 2018; Raja et al., 2011; Karegar et al., 2022; Zhang-Kennedy et al., 2016; Suh et al., 2022). For example, physical security metaphors can improve users’ understanding of personal firewalls (Raja et al., 2011). In other cases, however, metaphors have been less effective. For example, descriptions of end-to-end encryption using metaphors failed to improve understanding (Demjaha et al., 2018). Prior work has also begun to explore the effectiveness of metaphors specifically for explaining DP (Karegar et al., 2022). They find that functional metaphors can be useful for explaining both that injected randomness protects privacy and that there exists a tradeoff between privacy and accuracy. The metaphors we develop are also functional (i.e., focused on what DP offers), rather than structural (i.e. focused on how DP works) (Alaqra et al., [n. d.]). While the metaphors from prior work aim to cover a long list of facts about DP, they are not designed to emphasize the different kinds of disclosures against which DP may or may not protect—the focus of our work.
@@
Diagrams Another strategy for explaining PETs is the use of visualizations. For example, hypothetical outcome plots (Hullman et al., 2015) have been used to visualize the protection offered by DP (Smart et al., 2022; Pearce and Jiang, 2020); they have also been used to visualize DP’s accuracy implications for data curators (Nanayakkara et al., 2022). In the case of randomized response (Warner, 1965)—a simple instantiation of local DP—the injected noise can be represented through a spinner (Bullek et al., 2017; Dekel et al., 2023). Recent work has also explored the use of diagrams and animations in the specific context of location privacy (Xiong et al., 2022). Diagrams have also been used to explain other PETs such as encryption (Stransky et al., 2021). We build on this prior work to develop diagrams for DP in both the local and central models.
@@
Nutrition Labels One influential approach in privacy communication broadly has been the use of “nutrition labels” for privacy (Kelley et al., 2009). Drawing inspiration from standardized nutrition labels on food products, privacy labels have been proposed as an alternative or supplement to typical privacy policies with their notorious usability issues (Obar and Oeldorf-Hirsch, 2020; Turow et al., 2018). Privacy labels have proven to be a useful way to present privacy-related information (Kelley et al., 2009). Organizing key information into carefully-designed labels helps users find information more quickly than they would by perusing a traditional privacy policy (Kelley et al., 2010). Although originally proposed for websites, similar labels have since been developed for datasets (Holland et al., 2020) and Internet of Things devices (Emami-Naeini et al., 2022). Nutrition labels have even been proposed for describing DP (Xiong, 2020; Cummings et al., 2021). Apple has recently integrated the nutrition labels approach into their iOS app ecosystem. Unfortunately, the utility of these labels has been hampered by the fact that labels are not always easy to find and can be misleading or inaccurate (Cranor, 2022; Kollnig et al., 2022). Our work adapts the privacy label concept for the purpose of explaining DP—specifically for explaining how local and central DP may or may not protect against particular disclosure risks.
@@
Mental models refer to simplified versions of complex processes that people mentally hold and which may help them understand key pieces of information (Craik, 1967; Lin et al., 2012). Researchers have argued that mental models are important for effectively communicating security risks to end users (Stewart and Lacey, 2012). Camp (Camp, 2009) argues that a medical or public health mental model is particularly useful for conveying the implications of malicious code—in particular, “that everyone is at risk,” “the importance and continued autonomy in the face of risk” and the ”shared responsibility for community health.” In this way, mental models can rely on people’s existing knowledge to help them better grasp attributes of a new setting.
However, flawed mental models can lead to dangerous decisions (Vaniea et al., 2014; Wash, 2010). For example, Wash (Wash, 2010) proposes folk models of viruses and hackers and describes how these models help explain why people ignore security advice. A mental models approach can also clarify how people’s backgrounds may impact their understanding of risks (Kang et al., 2015; Oates et al., 2018; Bravo-Lillo et al., 2011). For instance, people’s level of computer science background impacts the complexity of their internet mental models, and therefore the number of privacy threats they perceive (Kang et al., 2015). Oates et al. (Oates et al., 2018) find that when asked to create illustrations of the meaning of privacy, experts’ illustrations tend to depict privacy as more “nuanced” than non-experts’ illustrations. Bravo-Lillo et al. (Bravo-Lillo et al., 2011) also find that novice and advanced users have different mental models and risk perceptions.
Finally, researchers have noted the value of studying privacy expectations (Lin et al., 2012; Rao et al., 2016). For example, Lin et al. (Lin et al., 2012) propose evaluating mobile app privacy by studying people’s privacy expectations of apps, while Rao et al. (Rao et al., 2016) suggest that understanding misalignment’s between people’s expectations and privacy policies can help reduce privacy risks.
3. Interview Study
@@
We situate our designs within the medical data collection scenario from (Cummings et al., 2021). In this scenario, a non-profit organization is collecting health data for medical research. Because medical information is considered highly sensitive (Ion et al., 2011; Schomakers et al., 2019), data subjects are more likely to care about understanding the privacy implications of DP in this scenario.
3.1. Initial Prototypes

### v2

@@ Abstract
1 Introduction
@@
Usable privacy research has long focused on making information about privacy protections transparent to end users in order to allow for informed decision-making about data sharing [64, 32, 30]. Prior work has sought to explain privacy enhancing technologies (PETs) such as end-to-end encryption in messaging [22] and HTTPS/TLS [33]. Existing explanation strategies vary in terms of how much they try to explain data protection mechanisms or processes versus implications of the particular PET. As PETs become increasingly complex and offer more nuanced notions of privacy, there is a significant need for increased research into best practice for transparent PET messaging that matches these new techniques and empowers end users.
@@
Differential privacy (DP) [28] is a privacy-enhancing technology that has quickly been adopted by industry and government agencies [31, 70, 97, 2, 72, 23]. DP deployments provide provable privacy guarantees by adding statistical noise to data or computations; this noise obfuscates the information of each individual while preserving aggregate-level insights. In response to DP’s rapid success, a growing body of work has started to document the inadequacies of existing messaging around DP [19] and design new messaging techniques for DP systems [105, 106, 51, 77, 13, 92, 34].
@@
The technical and mathematical complexity of PETs like DP makes effective communication challenging [19, 76, 3, 22]. DP is an interesting case study for constructing transparent PET messaging because it is an instance of an emerging PET paradigm that has received relatively little attention: privacy-preserving, outsourced computation. This paradigm is increasingly important as more users rely on computationally-weak mobile devices [83]. In response to this growing need, new approaches to privacy-preserving computational outsourcing are being actively developed, both in industry [6, 71] and academia [61]. Although there has been some work on people’s mental models of other PETs in this category [49] and creating transparent messaging for functional encryption in particular [5], further work is needed. DP is particularly interesting because of its widespread deployment and use of statistical methods (as opposed to encryption).
@@
In this work, we develop messaging for DP that highlights the threat models that are implicit in different approaches to deploying DP. These implicit threat models are critical for end users to understand before sharing their data, as the chosen threat model may not provide protections against the classes of attackers about which they are concerned. We do this by exploring three explanation formats drawn from the existing PETs messaging literature: nutrition labels [53], diagrams [106, 92, 94], and metaphors [108, 22, 84, 51, 109, 95]. Each of our evaluated explanations aims to communicate the consequences of DP in terms of which information flows the deployment protects against. Prior work has tended to focus on explaining individual PETs in isolation, and while we focus on DP, we discuss how our designs may be modified to account for the effects of multiple PETs deployed together.
@@
Differential Privacy Deployment Models. There are multiple deployment models for DP, each of which is associated with a particular threat model. The two most widely deployed models are the central model [29] and the local model [52]. [^1] Although there are many variations of DP [24], we focus on the central and local models due to their popularity.
@@
The central model assumes there exists a data curator who is trusted to see raw data from individuals; the adversary can only access released, aggregate results. The data curator collects data from individuals, performs statistical analyses on the collected dataset, and then injects statistical noise into the results before release. This process limits the ability of the adversary to reconstruct individual records from summary statistics, at the cost of reduced accuracy. The potential danger of this model, however, is that the data curator might not be trustworthy; the database storing individuals’ data could be vulnerable to hackers or misuse by insiders if other, complementary security practices are not adopted in tandem. Well-known deployments of the central model include the U.S. Census Bureau’s data products for the 2020 Decennial Census [2].
@@
In the local model, noise is added to each individual’s data before collection, meaning the unmodified data are never stored together. As such, there is not the same need to trust the data curator (i.e., it is assumed that the data curator is honest but curious). This higher level of security comes at a cost: significantly more noise must be added to the data in order to ensure the same level of privacy protection, reducing the accuracy—and, thus, utility—of the collected data. Notably, Google and Apple have both used local DP to analyze browser data in Chrome and Safari, respectively [31, 7].
@@
Data subjects cannot be expected to make informed data-sharing decisions if they believe that DP is “some sort of crypto-magic to protect people from data misuse” [89]. Prior work has demonstrated that existing DP description strategies do a poor job aligning users’ privacy expectations with the privacy protections provided by DP models [19]. In other words, the kind of protection that users expect does not align with the actual nature of the protection offered by DP. Misaligned expectations exacerbated by poor communication can lead to data subjects underestimating or—even more alarmingly—overestimating the privacy protection that DP offers.
@@
We first explore three kinds of explanations that build on prior work  [51, 78, 106, 53, 84, 19] through an interview study: metaphors, diagrams, and privacy labels for DP. Based on the results of this study, we identify the most promising strategies—privacy labels and metaphors—and further refine these explanations based on participant feedback. We evaluate our refined explanations in an online survey ($n=698$), measuring objective comprehension, subjective understanding, perceived thoroughness, and trust. We compare our explanations against existing state-of-the-art text-based explanations of DP [105]. Based on our results, we make suggestions for future research and design of explanations of PETs.
@@
2 Background
A growing body of work provides guidance for effective S&P communication [90, 36]. Awkward interfaces or ineffective communication can lead to dangerous misconceptions and risky behaviors [103, 59, 19, 39]. One reason that people may misjudge privacy risks or misuse PETs is that they lack appropriate mental models. Prior work has argued that “efficacy of risk communication depends not only on the nature of the risk, but also on the alignment between the conceptual model embedded in the risk communication and the user’s mental model of the risk” [8]. Unfortunately, existing depictions of DP appear to be misaligned with people’s mental models, resulting in misaligned privacy expectations [19].
@@ 2 Background
Table 1: Five Information Disclosures. We combine the “organization” and “data analyst” categories from prior work, since a data analyst is simply an employee of the organization [19]. Although some implementations of the central model limit employees’ access to the data (e.g., Uber [47]), we consider the more common case where only published information is privacy-protected.
@@
Implications vs Process. One of the most important findings from prior work is that explaining data protection processes is not enough for most readers to grasp the implications of the protection offered by PETs [92, 19, 105, 26]. Xiong et al. [105] studied explanations of both central and local DP, and found that when the implications of the local and central models were stated explicitly, participants were more willing to share information under the local model. Kühtreiber et al. [60] replicated this study with German participants. Differently from these studies, we explore a variety of best-practice methods from the usable S&P literature (i.e., metaphors, diagrams, and nutrition labels) to communicate which information flows are protected by DP under the central and local models.
@@
Cummings et al. [19] explored the implications of DP through six information disclosures about which people care and DP may protect—depending on whether the local or central model is used. They found that existing descriptions of DP fail to appropriately set privacy expectations regarding these disclosures, in part because many descriptions are not specific to the model (e.g., central or local) being used. In contrast to this work, we build new explanations rather than evaluating existing ones. We draw on their framework to present the implications of DP (entities to whom data can potentially be disclosed) as part of our nutrition labels. For improved clarity, in our study, we combine two of the disclosures (organization and data analyst), resulting in five total (Table 1).
@@
Finally, Frazen et al. [34] and Nanayakkara et al. [77] developed methods of explaining the implications of the privacy budget, drawing from the risk communication literature. Nanayakkara et al. [77] found that participants were more willing to share information as the privacy budget decreased (i.e., protections were strengthened). In our study, we assume a small privacy budget (i.e., strong privacy), so that we can focus on the implications of the deployment model. Future work could consider combining our explanations with explanations of the privacy budget.
@@
Metaphors Metaphors are one approach for improving mental models, and have been studied extensively in the S&P domain [108, 22, 84, 51, 109, 95]. For example, physical security metaphors can improve users’ understanding of personal firewalls [84]. In other cases, however, metaphors have been less effective. For example, descriptions of end-to-end encryption using metaphors failed to improve understanding [22]. Prior work has also begun to explore the effectiveness of metaphors specifically for explaining DP [51]. They find that functional metaphors can be useful for explaining both that injected randomness protects privacy and that there exists a tradeoff between privacy and accuracy. The metaphors we develop are also functional (i.e., focused on what DP offers), rather than structural (i.e. focused on how DP works) [4]. While the metaphors from prior work aim to cover a long list of facts about DP, they are not designed to emphasize the different kinds of disclosures against which DP may or may not protect—the focus of our work.
@@
Diagrams Another strategy for explaining PETs is the use of visualizations. For example, hypothetical outcome plots [44] have been used to visualize the protection offered by DP [92, 82]; they have also been used to visualize DP’s accuracy implications for data curators [75]. In the case of randomized response [101]—a simple instantiation of local DP—the injected noise can be represented through a spinner [13, 21]. Recent work has also explored the use of diagrams and animations in the specific context of location privacy [106]. Diagrams have also been used to explain other PETs such as encryption [94]. We build on this prior work to develop diagrams for DP in both the local and central models.
@@
Nutrition Labels One influential approach in privacy communication broadly has been the use of “nutrition labels” for privacy [53]. Drawing inspiration from standardized nutrition labels on food products, privacy labels have been proposed as an alternative or supplement to typical privacy policies with their notorious usability issues [81, 99]. Privacy labels have proven to be a useful way to present privacy-related information [53]. Organizing key information into carefully-designed labels helps users find information more quickly than they would by perusing a traditional privacy policy [54]. Although originally proposed for websites, similar labels have since been developed for datasets [43] and Internet of Things devices [30]. Nutrition labels have even been proposed for describing DP [104, 19]. Apple has recently integrated the nutrition labels approach into their iOS app ecosystem. Unfortunately, the utility of these labels has been hampered by the fact that labels are not always easy to find and can be misleading or inaccurate [18, 58]. Our work adapts the privacy label concept for the purpose of explaining DP—specifically for explaining how local and central DP may or may not protect against particular disclosure risks.
@@
Mental models refer to simplified versions of complex processes that people mentally hold and which may help them understand key pieces of information [17, 63]. Researchers have argued that mental models are important for effectively communicating security risks to end users [93]. Camp [14] argues that a medical or public health mental model is particularly useful for conveying the implications of malicious code—in particular, “that everyone is at risk,” “the importance and continued autonomy in the face of risk” and the ”shared responsibility for community health.” In this way, mental models can rely on people’s existing knowledge to help them better grasp attributes of a new setting.
However, flawed mental models can lead to dangerous decisions [100, 102]. For example, Wash [102] proposes folk models of viruses and hackers and describes how these models help explain why people ignore security advice. A mental models approach can also clarify how people’s backgrounds may impact their understanding of risks [50, 80, 12]. For instance, people’s level of computer science background impacts the complexity of their internet mental models, and therefore the number of privacy threats they perceive [50]. Oates et al. [80] find that when asked to create illustrations of the meaning of privacy, experts’ illustrations tend to depict privacy as more “nuanced” than non-experts’ illustrations. Bravo-Lillo et al. [12] also find that novice and advanced users have different mental models and risk perceptions.
Finally, researchers have noted the value of studying privacy expectations [63, 85]. For example, Lin et al. [63] propose evaluating mobile app privacy by studying people’s privacy expectations of apps, while Rao et al. [85] suggest that understanding misalignment’s between people’s expectations and privacy policies can help reduce privacy risks.
3 Interview Study
@@
We situate our designs within the medical data collection scenario from [19]. In this scenario, a non-profit organization is collecting health data for medical research. Because medical information is considered highly sensitive [45, 91], data subjects are more likely to care about understanding the privacy implications of DP in this scenario.
3.1 Initial Prototypes